### v1

# Learning Internal Biological Neuron Parameters and Complexity-Based Encoding for Improved Spiking Neural Networks Performance

Abstract
This study introduces a novel approach by replacing the traditional perceptron neuron model with a biologically inspired probabilistic meta neuron, where the internal neuron parameters are jointly learned, leading to improved classification accuracy of spiking neural networks (SNNs). To validate this innovation, we implement and compare two SNN architectures: one based on standard leaky integrate-and-fire (LIF) neurons and another utilizing the proposed probabilistic meta neuron model. As a second key contribution, we present a new biologically inspired classification framework that uniquely integrates SNNs with Lempel-Ziv complexity (LZC) a measure closely related to entropy rate. By combining the temporal precision and biological plausibility of SNNs with the capacity of LZC to capture structural regularity, the proposed approach enables efficient and interpretable classification of spatiotemporal neural data, an aspect not addressed in existing works. We consider learning algorithms such as backpropagation, spike-timing-dependent plasticity (STDP), and the Tempotron learning rule. To explore neural dynamics, we use Poisson processes to model neuronal spike trains, a well-established method for simulating the stochastic firing behavior of biological neurons. Our results reveal that depending on the training method, the classifier’s efficiency can improve by up to 11.00%, highlighting the advantage of learning additional neuron parameters beyond the traditional focus on weighted inputs alone.
keywords:
Spiking Neural Networks , Learning Algorithms , Neuron Models , Lempel-Ziv Complexity , Spike Trains, Classification
journal: Title
{highlights}
Joint neuron parameter and synaptic learning improves classification accuracy
SNN-LZC framework enables efficient, interpretable spatiotemporal classification
Validated across learning rules, the method shows robustness and flexibility
Supports biologically grounded models and rules for biosignal classification
Findings inform biologically inspired SNN design for classifications tasks
1 Introduction
Spiking Neural Networks (SNNs) have emerged as promising candidates for addressing the limitations of conventional artificial neural networks (ANNs), particularly in real-world applications where temporal precision and energy efficiency are critical [1]. Unlike traditional ANNs, SNNs more closely approximate the behavior of biological neurons by transmitting information through discrete spike events, thus capturing both the spatial and temporal dynamics of neural activity [3, 2]. This biologically grounded architecture positions SNNs as potential computational models capable of approaching the efficiency and adaptability of the human brain [4].
Despite their promise, training SNNs remains a significant challenge due to their non-linear, event-driven dynamics and the non-differentiable nature of spike generation [6, 7]. Traditional gradient-based methods such as backpropagation, which iteratively minimize a loss function by computing gradients of continuous activation values, are not well-suited to the discrete and temporally sparse nature of spiking activity [5]. Moreover, these conventional arithmetic-based learning rules are often biologically implausible and inefficient when applied to spiking architectures [8]. Instead, the optimization of intrinsic neuron parameters and the adoption of biologically inspired temporal coding schemes play a critical role in improving the performance and interpretability of SNNs. Temporal coding, in particular, has been shown to transmit information more efficiently than rate coding, potentially reducing the number of spikes required for accurate representation and decision-making [9].
On the other hand, concepts from Information Theory (IT), like Lempel-Ziv Complexty (LZC) provide powerful tools for analyzing neural data and understanding the computational principles underlying brain-like processing [10, 11, 12]. LZC offers a multidimensional perspective on neural signals [13, 14]. This quantity not only account for the probabilistic characteristics of spiking data but also capture the underlying structural patterns and temporal regularities. This dual sensitivity allows them to quantify the degree of randomness in neural activity while simultaneously identifying deterministic features, i.e. an essential property for characterizing the outputs of SNNs.
In this study, we propose a novel modification to the spiking neural network by replacing the traditional perceptron neuron model with a biologically inspired meta neuron, in which internal parameters are jointly learned during training. We hypothesize and demonstrate that this biologically motivated enhancement can significantly improve SNN classification accuracy. To validate this, we implement two SNN architectures: one based on standard leaky integrate-and-fire (LIF) neurons [15] and the other on the proposed meta neuron model [16]. As a second key contribution, we introduce a novel biologically inspired classification framework that integrates SNNs with Lempel-Ziv complexity. By leveraging the temporal precision and biological plausibility of SNNs together with LZC’s ability to quantify structural complexity, this framework provides efficient and interpretable analysis of spatiotemporal neural patterns for classification tasks, namely a capability not present in prior approaches. We further evaluate multiple learning algorithms, including backpropagation [5], spike-timing-dependent plasticity (STDP) [17], and the Tempotron learning rule [18, 19].
Notation Description
$\textbf{x}=[x_{1},x_{2},\cdots,x_{n}]\in\mathbb{R}^{n}$ synaptic weight vector
$V_{m}(t)\in\mathbb{R}$ membrane potential at time $t$
$\tau_{m}\in\mathbb{R}^{+}$ membrane time constant
$I(t)\in\mathbb{R}$ input current at time $t$
$\tau_{m}\in\mathbb{R}^{+}$ membrane time constant
$t_{f}\in\mathbb{R}^{+}$ firing time of neuron
$S_{i}(t)\in\mathbb{Z}$ binary-valued spike train (0 or 1 at any given time)
$t_{i}^{k}\in\mathbb{R}^{+}$ spike times (real-valued, positive)
$\eta\in\mathbb{R}^{+}$ general learning rate (positive real number) for weight update in supervised learning and gradient-based optimization
$\eta_{+},\eta_{-}\in\mathbb{R}^{+}$ STDP time constant (positive real)
$A_{-},A_{+}\in\mathbb{R}^{+}$ amplitude of STDP weight changes (positive real)
$t_{pre},t_{post}\in\mathbb{R}^{+}$ pre- and postsynaptic spike time (positive real)
$t_{spike}\in\mathbb{R}^{+}$ spike timing
$S_{pre}(t),S_{post}(t)\in{0,1}$ binary spike trains
$K(t)\in\mathbb{R}\longrightarrow\mathbb{R}$ kernel function, maps time to real value
$t_{i}^{target},t_{i}^{actual}\in\mathbb{R}^{+}$ targeted and ctual spike times
$\mathcal{E}\in\mathbb{R}^{+}$ error function (positive real)
$\frac{\partial E}{\partial w_{i}}\in\mathbb{R}$ gradient of error
Table 1: Basic notation used in the study.
2 Notation
An overview of the key notation is provided in Table 1.
Figure 1: Comparison of the standard Leaky Integrate-and-Fire (LIF) neuron model and the extended meta-neuron architecture. The LIF neuron integrates the input current $I(t)$ over time according to the leaky membrane equation $\tau\frac{dV}{dt}=-V+RI(t)$, where $\tau$ is the membrane time constant and $R$ is the membrane resistance. A spike is emitted when the membrane potential $V(t)$ exceeds a fixed threshold $V_{\text{th}}$, followed by a reset to $V_{\text{reset}}$. The meta-neuron preserves this core structure but incorporates dynamic, adaptive components. Specifically, the membrane time constant $\tau(t)$ and threshold $V_{\text{th}}(t)$ can vary over time, enabling richer and more flexible temporal dynamics. More generally, the meta-neuron can be interpreted as computing a parameterized function $y=f(\mathbf{x};\boldsymbol{\theta})$, where the input vector $\mathbf{x}$ is mapped to an output $y$ through parameters $\boldsymbol{\theta}$, which may be adapted online through feedback mechanisms or learning algorithms. This structure supports meta-learning and context-aware adaptation while maintaining biologically inspired spiking behavior.
3 Neuron Models
The Leaky Integrate-and-Fire (LIF) neuron receives an input vector $\mathbf{x}(t)\in\mathbb{R}^{n}$, which is linearly combined with a weight vector $\mathbf{w}\in\mathbb{R}^{n}$ and a bias term $b\in\mathbb{R}$ to compute the total input current:
$I(t)=\mathbf{w}^{\top}\mathbf{x}(t)+b.$
The membrane potential $V_{m}(t)$ evolves over time according to a leaky integration mechanism governed by the membrane time constant $\tau_{m}$ and resistance $R_{m}$. When $V_{m}(t)$ exceeds a fixed threshold $V_{\mathrm{th}}$, a spike is emitted and the potential is reset to the resting potential $V_{\mathrm{reset}}$ [15, 20].
In contrast, the meta-neuron extends this model by allowing key neuronal parameters to adapt over time [21]. Specifically, both the membrane time constant $\tau_{m}(t)$ and the threshold $V_{\mathrm{th}}(t)$ are dynamically modulated based on the internal state of the network or external feedback. More abstractly, the metaneuron computes an output.
$y(t)=f\!\big{(}\mathbf{x}(t);\boldsymbol{\theta}(t)\big{)},$
where $\boldsymbol{\theta}(t)$ denotes a time-varying set of internal parameters subject to learning or contextual updates. This framework enables more flexible and expressive computations while preserving the spiking behavior characteristic of biologically inspired models.
While the LIF model offers a biologically grounded mechanism for spike generation, its static parameters limit expressivity. To overcome this, we extend the LIF framework into a meta-neuron formulation, where parameters evolve dynamically and computations generalize beyond simple thresholding. As illustrated in Figure 1, the meta-neuron preserves the core integration-threshold-reset structure of the LIF but augments it with time-varying parameters such as $\tau(t)$ and $V_{\text{th}}(t)$, or even fully parameterized update functions. This supports a broader class of computations, bridging spiking dynamics with principles from meta-learning and optimization.
4 Spiking Neural Network
Figure 2: Schematic of the spiking neural network and its post-processing through Lempel–Ziv Complexity. The architecture includes an input layer $X=\{x^{(i)}\}$, hidden layer $H=\{h^{(i)}\}$, and output layer $Z=\{z^{(i)}\}$, each consisting of $n$ neurons. Binary input sequences of length 1024 are encoded into $n$-bit spike trains and propagated through the network. The output activity of each neuron $z^{(i)}$ is converted into a binary sequence $\mathbf{x}_{n}=[x_{1},x_{2},\dots,x_{n}]$, which is parsed into substrings to build a dictionary $\mathcal{D}$ of unique patterns. The LZC is computed as $c_{\alpha}(\mathbf{x}_{n})=\frac{C_{\alpha}(\mathbf{x}_{n})}{{n}}\log_{\alpha}n$, where $C_{\alpha}(\cdot)$ counts the number of distinct substrings that appear consecutively along the sequence. This measure quantifies the spatiotemporal complexity of the spike output.
The network architecture under consideration is made up of three distinct neuronal layers: input, hidden, and output. Each containing $n$ computational units, as illustrated in Figure 2. The input layer receives binary sequences of length 1024, which are encoded in spike-train representations of $n$-bit width and processed through the network. The spike trains produced by the output neurons $\{z^{(i)}\}$ are decoded into binary sequences $\mathbf{x}_{n}=[x_{1},x_{2},\dots,x_{n}]$, which are analyzed using the Lempel–Ziv Complexity metric. The LZC measures the number of unique substrings $C_{\alpha}(\mathbf{x}_{n})$ encountered during sequential parsing. To normalize this measure across sequence lengths, we compute $c_{\alpha}(\mathbf{x}_{n})=\frac{C_{\alpha}(\mathbf{x}_{n})}{{n}}\log_{\alpha}n$, where $\alpha=2$ for binary sequences [11]. This provides a compact descriptor of the structural complexity of the network’s output activity. This hybrid approach combines the temporal precision and biological plausibility of SNNs with the capacity of LZC to quantify the structural complexity of spike patterns. By leveraging these complementary properties, the model enables efficient and interpretable classification of spatiotemporal neural data [11, 13, 14]. This integration also enhances robustness to noise and improves performance, particularly when processing signals with variable temporal dynamics, such as those generated by Poisson processes [22, 23, 24, 25].
Poisson spike trains are produced using a homogeneous Poisson process, characterized by the independent occurrence of events at a fixed mean rate $\lambda$. The probability of observing a spike within a small time interval $\Delta t$ is given by $P(\lambda\Delta t)=o(\Delta t)$. The inter-spike intervals (ISIs) follow an exponential distribution $P(\lambda\Delta t)=\lambda e^{-\lambda\Delta t}$. Thus, spike sequences are generated with specific rates $\lambda_{x}$ and $\lambda_{y}$ for different input channels. To assess the influence of temporal variability on network dynamics and classification accuracy, these spike trains are evaluated under multiple rate conditions. The binary sequence of each process is converted into inputs to neuron models with specified $\Delta t=1ms$.
5 Learning Rules
Backpropagation (BP) is a gradient-based learning algorithm used to train neural networks by minimizing a loss function $\mathcal{E}$ [26]. The network parameters are updated through weight update
$\Delta w_{i}=-\eta\frac{\partial\mathcal{E}}{\partial w_{i}},\quad\frac{\partial\mathcal{E}}{\partial w_{i}}=\frac{\partial\mathcal{E}}{\partial V_{m}}\cdot\frac{\partial V_{m}}{\partial w_{i}},$
where $\eta$ is the learning rate and $V_{m}$ denotes an intermediate potential. BP propagates error gradients backward through the network to optimize weights [27], see Figure 3.
Figure 3: SNN trained with backpropagation learning. Output spike trains $\mathbf{z}^{(i)}$ are analyzed by LZC for classification. Error signals (red) propagate globally from output to input.
Figure 4: SNN trained using STDP. Local spike timing between pre- and post-synaptic neurons adjusts weights.
On the other hand, STDP is a biologically grounded, unsupervised learning mechanism in which synaptic weight changes depend on the precise relative timing of presynaptic and postsynaptic spikes [28], see Figure 4. When a presynaptic spike precedes a postsynaptic spike within a critical window, the synapse is potentiated (long-term potentiation, LTP); if the order is reversed, it undergoes depression (long-term depression, LTD). The weight update is defined as [29, 30, 31]:
$\Delta w_{i}=\begin{cases}A_{+}\exp\!\Big{(}-\frac{t_{\mathrm{post}}-t_{\mathrm{pre}}}{\tau_{+}}\Big{)},&t_{\mathrm{post}}>t_{\mathrm{pre}},\\[6.0pt] -A_{-}\exp\!\Big{(}-\frac{t_{\mathrm{pre}}-t_{\mathrm{post}}}{\tau_{-}}\Big{)},&t_{\mathrm{pre}}>t_{\mathrm{post}},\end{cases}$
where $A_{\pm}$ are amplitudes of potentiation/depression and $\tau_{\pm}$ are their time constants.
Figure 5: SNN trained using Tempotron rule. Local error signals flow from the output to hidden layer.
The Tempotron is a biologically inspired supervised learning rule designed for SNNs, which emphasizes temporal precision over spike rate [18]. At its core, it uses LIF neurons to integrate time-varying postsynaptic potentials (PSPs) from presynaptic spikes, see Figure 5. If the neuron should fire ($P^{+}$) or stay silent ($P^{-}$), weights are updated as:
$\Delta w_{i}=\eta\sum_{t\in T_{\mathrm{pre}}}K(t_{f}-t),$
where $T_{\mathrm{pre}}$ is the set of presynaptic spike times, $K(\cdot)$ is the postsynaptic potential kernel, $t_{f}$ is the time of maximal membrane potential, and $\eta$ is the learning rate. This corresponds to gradient descent on the peak membrane potential $V_{m}(t)$. This learning rule performs gradient descent on the peak membrane potential to minimize classification error. Compared to traditional rate-based approaches or STDP, the Tempotron requires fewer tunable parameters, operates under supervised feedback, and is optimally designed for temporal pattern recognition tasks where the timing of spikes conveys critical information.
6 Results
In our study, we introduce a novel spiking neural network framework that, to the best of our knowledge, is the first to integrate Lempel–Ziv Complexity with biologically inspired neuron models. We employ two distinct SNN architectures: one based on the widely used leaky integrate-and-fire neurons [15], and the other utilizing a biologically motivated meta neuron model [21]. This integration of LZC with SNNs provides a new classification approach, leveraging the temporal coding capabilities and neural plausibility of SNNs along with the quantification of structural complexity provided by LZC. The result is a robust and interpretable method for analyzing spatio-temporal data, an aspect that remains largely unexplored in the current SNN literature. To optimize performance and evaluate learning dynamics, we implement and compare several training algorithms, including backpropagation-based methods [5], spike-timing-dependent plasticity [17], and the Tempotron learning rule [18, 19].
Table 2 presents the classification accuracy results for spiking neural networks based on the leaky integrate-and-fire neuron model, using inputs generated as Poisson spike trains. Our findings indicate that optimizing learning and neuronal parameters can significantly improve classification accuracy. In particular, the application of Backpropagation and spike-timing-dependent plasticity led to accuracy improvements of up to 11.00% when both neuron parameters and the learning rate were optimized, as compared to the baseline configuration where only the learning rate was trained. In contrast, the Tempotron learning rule exhibited a more modest improvement of approximately 4.00%. However, its baseline performance, achieved by training only the learning rate, was already high, reaching 95.00% accuracy.
Table 2: Accuracy results for SNN based on LIF model.
Learning Parameters Backpropagation Tempotron STPD
Number of neurons
Learning rate
85.50% 95.00% 88.50%
Number of neurons
Learning rate
Thresholds
Decays
96.50% (11.00% $\uparrow$) 99.00% (4.00% $\uparrow$) 99.50%(11.00% $\uparrow$)
Table 3: Accuracy results for SNN based on meta neuron model.
Learning Parameters Backpropagation Tempotron STPD
Number of neurons
Learning rate
79.00% 85.00% 87.00%
Number of neurons
Learning rate
Thresholds
Decays
90.00% (11.00% $\uparrow$) 92.00% (7.00% $\uparrow$) 91.50%(4.50% $\uparrow$)
In Table 3 the classification accuracy results for spiking neural networks based on meta-neuron is shown. Similar to the LIF-based architecture, it was observed that jointly optimizing both the neuron parameters and the learning rate led to a noticeable improvement in classification accuracy. Among the evaluated learning algorithms, Backpropagation yielded the most substantial gain, with an increase of up to 11.00% relative to the baseline configuration where only the learning rate was adjusted.
Two optimization configurations were compared for each learning algorithm: (1) one can say baseline, where the number of neurons and the learning rate were optimized, and (2) extended, where intrinsic neuron parameters (thresholds and decay constants) were additionally optimized. For both the LIF and meta-neuron architectures, adding trainable thresholds and decay constants consistently improved classification accuracy. This suggests that adjusting firing thresholds enhanced the neuron’s sensitivity to relevant spike patterns, while tuning decay constants adapted the temporal integration window of post-synaptic potentials. The combined effect allowed the network to more effectively capture temporal dependencies in the input.
Backpropagation and STDP benefited the most from intrinsic parameter optimization, achieving up to 11% accuracy gain relative to the baseline. Tempotron, while already showing a strong baseline performance, exhibited smaller improvements (up to 4–7%), indicating that its performance relies primarily on synaptic weight dynamics rather than intrinsic parameter tuning. The smaller improvement observed for the Tempotron can be explained by the nature of its learning rule, which relies primarily on adjusting synaptic weights to shape the decision boundary. As a result, parameter tuning (thresholds and decay constants) contributes less to performance compared to Backpropagation or STDP. Thus,the optimization configurations and their interpretation was summarized in Table 4.
Table 4: Optimization configurations and their interpretation.
Configuration Optimized Parameters Interpretation and Effect
Baseline Number of neurons, learning rate Focuses on global capacity and convergence speed. Provides stable but limited adaptation to temporal patterns.
Extended Number of neurons, learning rate, thresholds, decay constants Adjusts firing sensitivity and temporal integration window. Increases responsiveness to relevant spike timings and improves accuracy, especially for Backpropagation and STDP.
7 Discussion
This study provides one of the first systematic comparisons of biologically inspired meta-neurons and classical LIF models in spiking neural networks, analyzing how their performance interacts with learning rules and input encoding strategies. By jointly exploring neuron dynamics, training algorithms, and Poisson-based input variability, our results offer novel insights into the trade-offs between expressiveness, stability, and accuracy in SNN classification tasks.
The LIF-based network consistently achieves above 96.50% accuracy, confirming its robustness under well-aligned input encodings [32, 33]. In contrast, the meta-neuron’s flexibility increases sensitivity to noisy Poisson input and hyperparameters, consistent with stability challenges discussed in [16]. Backpropagation improves meta-neuron performance by up to 11.00%, but does not close the gap to LIF accuracy. This aligns with the broader challenge of fully exploiting meta-neuron expressiveness without overfitting [34].
From a systems neuroscience perspective, the results highlight the potential of spiking neural networks not only as engineering tools, but also as computational models for probing fundamental questions of neural coding and learning. The observed sensitivity of classification performance to input stochasticity and spike timing mirrors variability found in biological systems [35], suggesting that SNNs can provide explanatory value for understanding sensory processing, neural variability, and adaptive behavior in biological circuits.
In practical terms, these insights can inform the design of neuromorphic architectures for low-power, real-time applications such as edge computing, autonomous systems, and brain–computer interfaces [32]. The complementary characteristics of stable LIF dynamics and adaptable meta-neuron components point toward hybrid architectures that balance interpretability with task-specific adaptability, i.e., well-suited for dynamic environments and non-stationary input streams. Furthermore, the strong dependence of expressive neuron models on precise hyperparameter tuning highlights the need for automated optimization pipelines. Approaches such as meta-learning and biologically inspired meta-plasticity [34] represent promising avenues to improve convergence, generalization, and adaptability without extensive manual calibration.
8 Conclusion
This study provides new insights on how the choice of learning algorithm, neuronal parameter optimization, and input encoding jointly shape the classification performance of spiking neural networks for temporally structured data. Although biologically inspired mechanisms such as STDP and the Tempotron rule offer valuable insights into synaptic plasticity, our results show for the first time that combining gradient-based methods such as backpropagation with flexible, parameter-optimized neuron models can yield significantly higher accuracy. Moreover, we also observed that under well-structured input datasets, simpler LIF-based architectures can outperform more complex models, revealing an important trade-off between biological plausibility, model complexity, and real-world performance.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Author contributions
All authors contributed to the conception and design of the study. Material preparation, data collection, and analysis were performed by all authors. The first draft of the manuscript was written by all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
Funding
Not applicable.
Data Availability Statement
Not applicable.